# Guiding Generative Models for Protein Design: Prompting, Steering and Aligning

### v1

@@ Abstract
Generative artificial intelligence (AI) models learn probability distributions from data and produce novel samples that capture the salient properties of their training sets. Proteins are particularly attractive for such approaches given their abundant data and the versatility of their representations, ranging from sequences to structures and functions. This versatility has motivated the rapid development of generative models for protein design, enabling the generation of functional proteins and enzymes with unprecedented success. However, because these models mirror their training distribution, they tend to sample from its most probable modes, while low-probability regions, often encoding valuable properties, remain underexplored. To address this challenge, recent work has focused on guiding generative models to produce proteins with user-specified properties, even when such properties are rare or absent from the original training distribution. In this review, we survey and categorize recent advances in conditioning generative models for protein design. We distinguish approaches that modify model parameters, such as reinforcement learning or supervised fine-tuning, from those that keep the model fixed, including conditional generation, retrieval-augmented strategies, Bayesian guidance, and tailored sampling methods. Together, these developments are beginning to enable the steering of generative models toward proteins with desired, and often previously inaccessible, properties.
@@ 1 Introduction
Learning to generate samples from complex probability distributions lies at the core of modern generative modeling. In the context of proteins, the availability of large datasets has catalyzed the rapid development of powerful generative models for protein design (GMPDs). Among GMPDs, some of the most adopted architectures are diffusion models, which reconstruct atomic coordinates by reversing a noise-adding process (Fig. 1A)[1, 2, 3], and protein language models (pLMs), trained either on masked sequence reconstruction (e.g. ESM[4] or the MSA Transformer[5]) or on autoregressive next-token prediction (e.g. ProtGPT2[6], ProGen2-3[7], and Evo-1/2[8, 9]) (Fig. 1B). In recent years, such models have enabled remarkable achievements, including the design of binders [10], ligand-binding receptors [11], and de novo enzymes [12] on unprecedented timescales.
@@
Figure 1: Generative modeling approaches and schematic illustrations of protein fitness landscapes. (A) Diffusion-based generative models. Protein structures are represented by the distribution $p(x)$ of atomic coordinates in three-dimensional space. During training, true atomic coordinates $X_{0}$ are progressively corrupted by Gaussian noise until reaching a normal distribution $\mathcal{N}(0,1)$. At inference, the model reverses this process, iteratively denoising random noise to reconstruct a protein structure. (B) Sequence-based generative models. Here, $p(x)$ denotes the distribution of amino acids across sequence positions. Training can proceed via masked-token prediction, in which the model infers the identity of masked residues, or via autoregressive next-token prediction, where sequences are generated one residue at a time. (C) Protein fitness landscape. Schematic illustration of the relationship between the data distribution $p(x)$ (blue points), typically learned in an unsupervised manner, and the conditional distribution $p(y|x)$ (red points), which focuses sampling on regions associated with high fitness. (D) Evolutionary accessibility. Evolution explores the fitness landscape through local, incremental mutations that can reach only contiguous high-fitness regions (solid paths), whereas transitions across fitness valleys (dashed paths) are inaccessible. In contrast, generative protein design models (GMPDs) can, in principle, traverse the landscape in a less constrained manner, directly sampling from otherwise evolutionarily inaccessible regions.
@@
Trained in a large corpus of natural sequences, structures, and functional annotations [e.g., 13, 14, 7, 15], GMPDs can explore vast regions of the protein sequence–structure landscape [e.g., 7, 16, 17, 18]. However, protein engineering often targets exceptional properties that are rare or even disfavored by natural selection (such as extreme thermostability in mesophilic organisms). These desirable functional optima may correspond to isolated high-fitness peaks separated by deep valleys in the evolutionary landscape (Fig. 1C, D), representing trajectories that evolution is unlikely to traverse. Consequently, GMPDs trained solely on natural data tend to assign negligible probability to these regions, making them difficult to sample directly. Further biases arise from uneven phylogenetic representation and residual annotation errors in the underlying datasets [e.g., 19, 20, 21, 22], which may further distort the learned landscape. In this sense, recent work highlight that simply increasing model scale does not guarantee monotonic improvements in fitness prediction [23].
@@
However, when appropriately guided, GMPDs have the potential to access low-probability yet functionally optimal regions of protein space. More formally, the protein design objective can be defined as generating protein sequences $x$ with desired properties $y$, where $y$ may represent attributes such as thermostability, catalytic efficiency, binding specificity, or other design goals. By contrast, the primary training objective of a GMPD is to model the distribution of natural proteins, $p(x)$. To bridge this gap, recent approaches aim to model the conditional distribution $p(x|y)$, thereby biasing generation toward sequences with the desired attributes (Fig. 1C). In many cases, this objective can also be viewed as an optimization problem, where the goal is to find $x$ that maximizes $p(x|y)$. Several recent techniques can be conceptually unified under this $p(x|y)$ framework, showing strong success in aligning GMPDs with specific design objectives. In this review, we categorize these strategies into two broad classes: (i) train-time methods, which modify the model parameters so that the learned distribution $p(x)$ approximates $p(x|y)$, and (ii) inference-time control, which guides generation toward target properties without altering the model’s underlying weights.
1.1 Train-time methods for controlled Generation
@@
Train-time methods modify the underlying probability distribution by directly updating GMPDs’ parameters. The most straightforward approach in this category is Supervised Fine-Tuning (SFT). In SFT, a pre-trained GMPD is further optimized under the same objective used during pre-training, but on a curated dataset of high-quality examples. For instance, a pre-trained protein language model (pLM) can be fine-tuned on a carefully assembled dataset from a target enzyme family to generate novel and distant members of that family. This process adapts the model’s parameters to a specific domain, effectively shifting its generative prior to align with the target data. SFT has achieved notable success, including the generation of enzymes [24, 25], gene editors[26, 27], or bacteriophages [28].
Method Description Examples
@@
Train-time methods Supervised finetuning (SFT) Fit the model to well curated data, shifting the learned distribution towards the data.
@@
Reinforcement Learning Aligns the model on feedback data over model’s outputs. Methods include preference- or reward-driven learning. PPO, GRPO, DPO
@@
Inference-time control Prompt & context programming Generation guided by structuring the input prompt with explicit instructions or templates (e.g., for specific positions or motifs). Masked pLM, EvoDiff, BoltzGen
@@
Retrieval-Augmented Generation (RAG) Enhances generation by dynamically incorporating external knowledge retrieved from a large corpus. RAG pipelines
Output-dependent guidance Gradient-based guided generation in sequence space based on the inference output. PPLM, ColabDesign, BoltzDesign, BindCraft
@@
Activation steering Direct manipulation of hidden states (e.g., residual stream) to promote or suppress attributes without parameter updates. Sparse Autoencoders (SAEs), steering vectors
@@
Bayesian guidance Re-weights the probability distribution using Bayes’ theorem: $\tilde{p}(y\mid x)\propto p_{\theta}(y\mid x)\exp(\lambda s(y,x))$. Bayesian scoring functions
@@
Sampling controls Alters sampling strategy (temperature, top-k, top-p, or more advanced searches like beam search and MCTS) to influence randomness and diversity. MCTS, temperature scaling, top-k, top-p
Table 1: Overview of Train-time methods and Inference-time Control Methods with Examples. pLM: protein language model. MCTS: Monte Carlo Tree Search.
@@
While SFT effectively specializes GMPDs toward generating samples representative of a particular dataset, it does not provide the model with the ability to discriminate by data quality, i.e., to differentiate among varying degrees of a desired property. Similar limitations have been observed in Natural Language Processing (NLP), where SFT alone often leads to suboptimal alignment with user intent and [29], in some cases, catastrophic forgetting [30]. Consequently, SFT is now commonly combined with reinforcement learning (RL) to achieve finer control over model behavior [e.g 31].
@@
In RL, a model learns to make optimal decisions by interacting with an environment through trial and error, receiving feedback in the form of rewards or penalties to maximize its cumulative reward over time. Unlike SFT, in RL the model is not provided with explicit examples of the desired outputs. Instead, the GMPD must infer and explore autonomously, potentially uncovering novel solutions that might not have been anticipated. More technically, a pre-trained model is treated as a policy $\pi_{\theta}(x)$ and updated to maximize a scalar reward -or equivalently, the probability over preferences- while constraining excessive deviation from its pre-trained distribution. Similarly to SFT, this process enables a transition from the unconditional distribution $p(x)$ toward the desired conditional distribution $p(x|y)$. Today, RL is central to the alignment of large language models (LLMs) and has driven remarkable advances across diverse fields, from autonomous driving to game playing.
@@
Several RL techniques have been implemented over the years. REINVENT (2017) [32] provides an early attempt, in the molecular field, to leverage vanilla RL policy-gradient methods (i.e., REINFORCE) to move from a broad distribution $p(x)$ toward a desired conditional distribution $p(x|y)$. REINVENT re-frames the popular REINFORCE update as an "augmented-likelihood objective". This objective allows to increase a property $y$’s score, while retaining probability mass near the original distribution $p(x)$, allowing the steering samples into high-scoring regions without sacrificing realism. Shortly after, Proximal Policy Approximation (PPO, 2017) [33] was introduced, providing a key improvement over vanilla policy-gradient methods; its clipped surrogate objective approximates a trust region, guaranteeing more stable updates.
Figure 2: Different stages where intervention can occur.
@@
PPO became the basis for one of the two main families of RL from Human Feedback (RLFH, 2022) approaches, allowing to align LLM to human preferences. In the first approach, a reward model is first trained on (human) ranked responses (i.e., preferences) for desired properties $y$ (e.g., clarity of text, non-offensive text). Subsequently, the reward model is used within the PPO update to align a pre-trained LLM, encoding $p(x)$, towards the properties $y$. Similarly to REINVENT, RLHF augments the reward with a penalty for drifting too far from the original distribution $p(x)$. The more recent GRPO algorithm [34] (2024) computes the rewards for groups of LLM’s responses, augmenting the PPO update with a reward baseline (i.e., computed across the group), without needing to train any additional value model. Direct Preference Optimization [35] (DPO, 2023) introduced the second family of RLHF approaches. DPO exploits the same preference signal as PPO-based approaches, but recasts alignment as a supervised objective on the log-probability differences between ranked (human) preferences. In practice, this dispenses the need to train an explicit reward model, provides a simpler and, potentially, more stable route toward preference alignment.
@@
These techniques have quickly met considerable success in the protein research realm, as summarized in table 2. PPO-style functional alignment has been used to update pLMs with experimental measurements, e.g., RL from eXperimental Feedback (RLXF) aligns an ESM-based generator toward brighter CreiLOV variants while constraining drift from the original distribution[36]. Model-based RL in the AlphaZero style has been applied to backbone construction, where Monte-Carlo Tree Search guided by a policy-value network outperforms plain tree search on top-down design tasks[37]. Preference-based tuning with DPO has been used to bias structure-conditioned models toward stability (preferring stabilizing over destabilizing sequences given a target backbone) and to reduce MHC-I epitope load while maintaining the fold[38, 39]. In parallel, mutation-policy RL frameworks propose on-policy sequence edits under specific oracles[36], and ProtRL extends the broad application of DPO and GRPO to protein engineering with pLMs, a framework that led to the design of low-nanomolar EGFR inhibitors[29].
@@
1.2 Inference-time control
In contrast to train-time methods, inference-time control methods guide GMPDs without updating their parameters, by directly influencing the models at the time of generation. These approaches assume that a pretrained GMPD already captures a rich representation of protein sequence-structure relationships, and that new behaviors can be elicited by steering the model’s inference process rather than retraining it. Such methods offer advantages in efficiency and flexibility, as the same model can be repurposed for diverse design tasks, enabling rapid exploration of new hypotheses while avoiding costly retraining.
Interventions can be applied at distinct stages of the generative process, from input to final sampling. Figure 2 summarizes these methodologies, categorized according to the site of intervention. Prompt and context engineering refers to shaping model behavior by modifying its inputs or conditioning information. pLMs can be trained with various control tags, such as Enzyme Commission numbers (ZymCTRL, [24]), Uniprot functional keywords (Progen, [25]), taxonomy (Evo1/2, [8, 9]), or combinations thereof [50]. Upon prompting these labels, the model specifically generates sequences for those cases. In other cases, specific residues may be masked and the model asked to reconstruct them, thereby exploring local sequence neighborhoods, or the model can be conditioned on predefined motifs and catalytic residues to guide functional design [51, 52, 53, 54, 10, 55]. Related to this idea, retrieval-augmented generation (RAG) has been introduced to dynamically incorporate external knowledge into the GMPD’s context[56, 57, 58]. By drawing upon semantically related examples from large databases, RAG enables more informed sampling and can enrich the design process with functional or structural priors not explicitly encoded in the pretrained model. A recent example is Protriever[57], which introduces a retrieval-augmented protein language model that jointly learns to retrieve homologous sequences and model their fitness, integrating evolutionary context at inference time without explicit structural supervision. Interventions can be applied at distinct stages of the generative process, from input to final sampling. Figure 2 summarizes these methodologies, categorized according to the site of intervention. Prompt and context engineering refers to shaping model behavior by modifying its inputs or conditioning information. pLMs can be trained with various control tags, such as Enzyme Commission numbers (ZymCTRL, [24]), Uniprot functional keywords (Progen1, [25]), taxonomy (Evo1/2, [8, 9]), or combinations thereof [50]. Upon prompting these labels, the model specifically generates sequences for those cases. In other cases, specific residues may be masked and the model asked to reconstruct them, thereby exploring local sequence neighborhoods[59], or the model can be conditioned on predefined motifs and catalytic residues to guide functional design [51, 60, 53]. Other approaches act directly on the latent encodings of protein sequences in a feedback-optimization loop, exploring regions of protein-structure space that satisfy user-defined design objectives such as ColabDesign, BindCraft and following methods [10, 54]. Alternatively, architectures such as BoltzGen[3] implement a conditional generative diffusion model with continuous guidance, in which encoded design conditions—such as binding-site specifications or structural constraints—are propagated throughout the denoising process to steer generation toward conformations consistent with the imposed design criteria [61] .

### v2

# Steering Generative Models for Protein Design: Aligning and Conditioning Strategies

@@ Abstract
Generative artificial intelligence models learn probability distributions from data and produce novel samples that capture the salient properties of their training sets. Proteins are particularly attractive for such approaches given their abundant data and the versatility of their representations, ranging from sequences to structures and functions. This versatility has motivated the rapid development of generative models for protein design, enabling the generation of functional proteins and enzymes with unprecedented success. However, because these models mirror their training distribution, they tend to sample from its most probable modes, while low-probability regions, often encoding valuable properties, remain underexplored. To address this challenge, recent work has proposed strategies for steering generative models toward user-specified properties. In this review, we survey and categorize these strategies, distinguishing approaches that modify model parameters, such as reinforcement learning or supervised fine-tuning, from those that keep the model’s parameters fixed, including conditional generation, retrieval-augmented strategies, Bayesian guidance, and tailored sampling methods. Together, these developments are beginning to enable the steering of generative models toward proteins with desired properties.
@@ 1 Introduction
Learning to generate samples from complex probability distributions lies at the core of modern generative modeling. In the context of proteins, the availability of large datasets has catalyzed the rapid development of powerful generative models (GM) for protein design. Some of the most adopted architectures are diffusion models, which reconstruct atomic coordinates by reversing a noise-adding process (Fig. 1A)[1, 2, 3], and protein language models (pLMs), trained either on masked sequence reconstruction (e.g. ESM [4] or the MSA Transformer[5]) or on autoregressive next-token prediction (e.g. ProtGPT2[6], ProGen2-3[7], and Evo-1/2[8, 9]) (Fig. 1B). In recent years, such models have enabled remarkable achievements, including the design of binders [10], ligand-binding receptors [11], and de novo enzymes [12] on unprecedented timescales. For broader overviews of generative AI for protein design, we refer the reader to recent reviews [13, 14, 15].
@@
Figure 1: Generative modeling approaches and schematic illustrations of protein fitness landscapes. (A) Diffusion-based generative models. Protein structures are represented by the distribution $p(x)$ of atomic coordinates in three-dimensional space. During training, true atomic coordinates $X_{0}$ are progressively corrupted by Gaussian noise until reaching a normal distribution $\mathcal{N}(0,1)$. At inference, the model reverses this process, iteratively denoising random noise to reconstruct a protein structure. (B) Sequence-based generative models. Here, $p(x)$ denotes the distribution of amino acids across sequence positions. Training can proceed via masked-token prediction, in which the model infers the identity of masked residues, or via autoregressive next-token prediction, where sequences are generated one residue at a time. (C) Protein fitness landscape. Schematic illustration of the relationship between the data distribution $p(x)$ (blue points), typically learned in an unsupervised manner, and the conditional distribution $p(y|x)$ (red points), which focuses sampling on regions associated with high fitness. (D) Evolutionary accessibility. Evolution explores the fitness landscape through local, incremental mutations that can reach only contiguous high-fitness regions (solid paths), whereas transitions across fitness valleys (dashed paths) are inaccessible. In contrast, generative protein design models can, in principle, traverse the landscape in a less constrained manner, directly sampling from otherwise evolutionarily inaccessible regions.
@@
Trained in a large corpus of natural sequences, structures, and functional annotations [e.g., 16, 17, 7, 18], GMs can explore vast regions of the protein sequence–structure landscape [e.g., 7, 19, 20, 21]. However, protein engineering often targets exceptional properties that are rare or even disfavored by natural selection (such as extreme thermostability in mesophilic organisms). These desirable functional optima may correspond to isolated high-fitness peaks separated by deep valleys in the evolutionary landscape (Fig. 1C, D), representing trajectories that evolution is unlikely to traverse. Consequently, GMs trained solely on natural data tend to assign negligible probability to these regions, making them difficult to sample directly. Further biases arise from uneven phylogenetic representation and residual annotation errors in the underlying datasets [e.g., 22, 23, 24, 25], which may further distort the learned landscape. In this sense, recent work highlight that simply increasing model scale does not guarantee monotonic improvements in fitness prediction [26].
@@
However, when appropriately guided, GMs have the potential to access low-probability yet functionally optimal regions of protein space. More formally, the protein design objective can be defined as generating protein sequences $x$ with desired properties $y$, where $y$ may represent attributes such as thermostability, catalytic efficiency, binding specificity, or other design goals. By contrast, the primary training objective of a GM is to model the distribution of natural proteins, $p(x)$. To bridge this gap, recent approaches aim to model the conditional distribution $p(x|y)$, thereby biasing generation toward sequences with the desired attributes (Fig. 1C). In many cases, this objective can also be viewed as an optimization problem, where the goal is to find $x$ that maximizes $p(x|y)$. Several recent techniques can be conceptually unified under this $p(x|y)$ framework, showing strong success in aligning GMs with specific design objectives. In this review, we categorize these strategies into two broad classes: (i) parameter-updating alignment, which modify the model parameters so that the learned distribution $p(x)$ approximates $p(x|y)$, and (ii) parameter-fixed steering, which guides generation toward target properties without altering the model’s underlying weights.
1.1 Parameter-Updating Alignment Modifies weights to shift $p(x)$
@@
Parameter-updating alignment modify the underlying probability distribution by directly updating GMs’ parameters. The most straightforward approach in this category is Supervised Fine-Tuning (SFT). In SFT, a pre-trained GM is further optimized under the same objective used during pre-training, but on a curated dataset of high-quality examples. For instance, a pre-trained protein language model (pLM) can be fine-tuned on a carefully assembled dataset from a target enzyme family to generate novel and distant members of that family. This process adapts the model’s parameters to a specific domain, effectively shifting its generative prior to align with the target data. SFT has achieved notable success, including the generation of enzymes [27, 28], gene editors[29, 30], or bacteriophages [31].
Category Method Intervention Local Description Examples
@@
Parameter-Updating Alignment Supervised finetuning (SFT) Model Weights Fit the model to well-curated data, shifting the learned distribution towards the data. [27, 28, 29, 30]
@@
Reinforcement Learning Aligns the model on feedback data over model’s outputs (preference or reward-driven learning). [32, 33, 34, 21]
@@
Parameter-Fixed Steering Prompt & context programming Prompting Generation guided by structuring the input prompt with explicit instructions or templates. [35, 36, 37, 38]
@@
Retrieval-Augmented Generation (RAG) Prompting Enhances generation by dynamically incorporating external knowledge from a large corpus. [39, 40, 41]
@@
Activation steering Internal Representation Direct manipulation of hidden states (e.g., residual stream) to promote or suppress attributes. [42, 43, 44]
Output-dependent guidance Decoding Gradient-based guided generation in sequence space based on the inference output. [10, 45]
@@
Bayesian guidance Decoding Re-weights the probability distribution using Bayes’ theorem: $\tilde{p}(y\mid x)\propto p_{\theta}(y\mid x)\exp(\lambda s(y,x))$. [46]
@@
Sampling controls Sampling Alters sampling strategy (temperature, top-k, top-p, MCTS) to influence randomness. [9, 47, 48]
Table 1: Overview of alignment and steering methods categorized by intervention stage within the generative process.
@@
While SFT effectively specializes GMs toward generating samples representative of a particular dataset, it does not provide the model with the ability to discriminate by data quality, i.e., to differentiate among varying degrees of a desired property. Similar limitations have been observed in Natural Language Processing (NLP), where SFT alone often leads to suboptimal alignment with user intent and [33], in some cases, catastrophic forgetting [49]. Consequently, SFT is now commonly combined with reinforcement learning (RL) to achieve finer control over model behavior [e.g 50].
@@
In RL, a model learns to make optimal decisions by interacting with an environment through trial and error, receiving feedback in the form of rewards or penalties to maximize its cumulative reward over time. Unlike SFT, in RL the model is not provided with explicit examples of the desired outputs. Instead, the GM must infer and explore autonomously, potentially uncovering novel solutions that might not have been anticipated. More technically, a pre-trained model is treated as a policy $\pi_{\theta}(x)$ and updated to maximize a scalar reward, or the probability over preferences, while constraining excessive deviation from its pre-trained distribution. Similarly to SFT, this process enables a transition from the unconditional distribution $p(x)$ toward the desired conditional distribution $p(x|y)$. Today, RL is central to the alignment of large language models (LLMs) and has driven remarkable advances across diverse fields, from autonomous driving to game playing.
@@
While different RL algorithms for model alignment have been proposed over the years, we can broadly group them in two main categories, 1) deep RL-based approaches and 2) direct preference learning. In deep RL-based approaches, a reward model is first trained on scored responses based on desired properties $y$ (e.g., human preferences). Subsequently, the reward model is used to update the policy (e.g., a pre-trained LLM), typically via a policy gradient update, shifting the underlying policy distribution $p(x)$ towards the properties $y$. REINVENT (2017) [51] provides an early attempt to apply deep RL-based approaches to the molecular field. REINVENT reframed the vanilla policy-gradient update (i.e., REINFORCE) as an "augmented likelihood objective". This objective enables adjusting the generative model parameters to favor samples with high scores on a target property $y$, while retaining probability mass near the original distribution $p(x)$, thereby preserving realism. Shortly after, Proximal Policy Approximation (PPO, 2017) [52] was introduced in the broader RL field, providing a key improvement over vanilla policy-gradient methods; its clipped surrogate objective approximates a trust region, guaranteeing more stable policy updates.
Figure 2: Different stages at which intervention can occur in protein generative modeling. Interventions may act on the model weights (Parameter-Updating Alignment), or without changing the base model parameters (Parameter-Fixed Steering). The latter category is further subdivided based on the stage of the model’s sampling pipeline where steering is applied: (1) specification of the input or context (e.g., prompting or retrieved context), (2) transformation into internal latent/activation representations, (3) decoding of latent representations into outputs, and (4) the sampling procedure that generates the final sequence or structure.
@@
The PPO update became the basis for RL from Human Feedback (RLHF) approaches [53], allowing to align LLM to human preferences, while including a penalty for when the aligned model drifts too far from the original distribution $p(x)$. The more recent GRPO algorithm [54] (2024) provides an alternative, where the reward is computed for a group of LLM’s responses, augmenting the PPO update with a reward baseline (i.e., computed across the group), without needing to train any additional value model. Direct Preference Optimization [55] (DPO, 2024) introduced the second category of RL approaches for model alignment: direct preference learning. DPO leverages the same preference signal as PPO-based methods, but cast alignment as a supervised objective over log-probability differences between ranked (human) preferences. In practice, this removes the need to train an explicit reward model, offering a simpler and potentially more stable route to preference alignment. These characteristics made DPO rapidly popular, and it has since spawned extensive variants of the original formulation.
@@
These techniques have quickly met considerable success in the protein research realm, as summarized in table 2. Deep RL-based approaches have been used to update pLMs with experimental measurements; for example, RL from eXperimental Feedback (RLXF) aligns an ESM-based generator toward brighter CreiLOV variants while constraining drift from the original distribution[56]. In another example, Monte-Carlo Tree Search guided by a policy-value network (AlphaZero-like [57]) has been applied to backbone construction, outperforming plain tree search on top-down design tasks[58]. In parallel, direct preference learning has been used to bias structure-conditioned models toward stability (preferring stabilizing over destabilizing sequences given a target backbone), to improve output quality within specific protein families, and to reduce MHC-I epitope load while preserving the fold [32, 21, 59]. Finally, recent frameworks aim to unify and extend these ideas: mutation-policy RL methods propose on-policy sequence edits under task-specific oracles [56], while ProtRL generalizes the application of REINVENT, DPO and GRPO algorithms to protein engineering with pLMs, exemplified on the design of low-nanomolar EGFR inhibitors [33].
While Direct preference learning approaches tend to be easier to work with compared to deep RL-based methods, it is still not clear which set of approaches lead to better performance under which conditions and, extensive benchmarking is needed [60], especially, in relation to protein design applications. Importantly, we believe the biggest limiting factor to the success of these methods in protein design is the difficulty of selecting good scoring metrics for most properties of interest.
@@
1.2 Parameter-fixed steering
While parameter-updating alignment methods form a relatively well-defined class of methods, we use the term parameter-fixed steering to refer to a broader and more heterogeneous family of approaches. Despite their diversity, these methods share a common objective: conditioning model generation toward a desired property $y$ without modifying the generative model’s base parameters. Conceptually, generation of learned GM can be decomposed into four stages: (1) an input or context specification stage (prompting), (2) the transformation of this input into internal latent or activation representations, and (3) the generation of the final output through the model’s decoding or (4) sampling procedure. The behavior of a GM can be influenced by intervening at any of these stages, as summarized in figure 2.
Intervention at the input or context specification stage (prompting)
. In the first stage, input conditioning specifies desired attributes through context or control tokens. When such attributes are provided during training, the model effectively learns conditional generation (sampling from an implicit $p_{\theta}(x|y))$ rather than post-hoc steering of an unconditional $p_{\theta}(x)$. For example, several pLMs have been trained with various control tags, such as Enzyme Commission numbers (ZymCTRL,[27]), Uniprot functional keywords (Progen1, [28]), taxonomy (Evo1/2, [8, 9]), or combinations thereof [70].
If the conditioning signal is not explicitly provided during training, desired constraints can instead be imposed at inference time via partial specification or hard constraints. For example, structure-conditioned sequence design models can fix selected residues (e.g., functional motifs) while redesigning the remaining positions [36], and atomic-context conditioning can bias sequence design in the presence of ligands, nucleotides, or metals [37]. More explicitly, some sequence generative models support infilling/inpainting, in which certain sequence segments are held fixed and the remaining residues are generated by masking and decoding the missing positions [35]. Others initialize sequences with unknown tokens (e.g., an all-X sequence) to induce structure hallucination in diffusion-based predictors, followed by iterative sequence redesign and structure re-prediction [38].
Other approaches act directly on the latent encodings of protein sequences in a feedback-optimization loop, exploring regions of protein-structure space that satisfy user-defined design objectives such as ColabDesign, BindCraft and following methods [10, 45]. It is important to note that in machine learning "generative" strictly refers to models that sample from a learned probability distribution $p(x)$. By this definition, techniques based on activation maximization, where a predictor is coupled with gradient-based optimization to produce sequences (e.g., BindCraft), are not generative in the formal sense, as they do not define or sample from a prior. However, in the protein design community, these methods are often informally described as generative and, here, we include such approaches due to their prominent usage and success.